### v1

# AI-PACE: A Framework for Integrating AI into Medical Education

Abstract
The integration of artificial intelligence (AI) into healthcare is accelerating, yet medical education has not kept pace with these technological advancements. This paper synthesizes current knowledge on AI in medical education through a comprehensive analysis of the literature, identifying key competencies, curricular approaches, and implementation strategies. The aim is highlighting the critical need for structured AI education across the medical learning continuum and offer a framework for curriculum development. The findings presented suggest that effective AI education requires longitudinal integration throughout medical training, interdisciplinary collaboration, and balanced attention to both technical fundamentals and clinical applications. This paper serves as a foundation for medical educators seeking to prepare future physicians for an AI-enhanced healthcare environment.
Keywords Medical Education $\cdot$ Artificial Intelligence $\cdot$ Curriculum Development $\cdot$ AI-PACE $\cdot$ Implementation Strategy
1 Introduction
Artificial intelligence’s (AI) role in medicine is rapidly changing; it is now an active clinical participant, from diagnostic imaging to administrative workflow optimization compared to an occasional aide in the past. This educational urgency is amplified by how rapidly consumers have adopted generative AI (GenAI). ChatGPT set the record for the fastest-growing user base, eclipsing previous record holders like TikTok and Instagram (Hu, 2023). With GenAI, patients are not waiting, they are forging ahead (Salmi, 2026; Stokel-Walker, 2024). They are already querying current models from OpenAI, Google, and Anthropic about medical conditions.
OpenAI stated that in January of 2026, 230 million customers were asking ChatGPT about health and wellness per week (OpenAI, 2026). Seeing market demands, There are now two health focused GenAI tools for from the frontier labs, Claude for Healthcare (Anthropic, 2026) and ChatGPT Health (OpenAI, 2026). It is now possible for physicians to encounter patients who could activate AI voice assistants mid-visit to interpret recommendations in real-time. Hence, we are facing a "shadow" health system that clinicians must be prepared to navigate, whether or not formal training has equipped them to do so.
Yet, current educational models fail to bridge the gap between the information age’s requirement for knowledge curation and the medical profession’s enduring need for deep biomedical competence (Wartman, 2019; Buja, 2019). While 44% of physicians report feeling inadequately prepared for new technologies (Minor, 2020), educational institutions have struggled to move beyond elective courses and short-term pilots. This disconnect presents a critical risk: a workforce expected to partner with AI tools without the foundational literacy to trust, verify, or oversee them.
Current educational efforts are often fragmented or siloed within specific domains. As evidenced by the short-term successes in specialty-specific pilots (van Kooten and others, 2024), such concentrated interventions risk rapid knowledge decay without longitudinal reinforcement. Similarly, Singla et al. (Singla and others, 2024) established a robust consensus on what competencies undergraduate medical students need but did not prescribe how these should evolve across the continuum of training. Other frameworks remain specialty-specific, such as Valikodath (Valikodath and others, 2021) defined for Ophthalmology.
This leaves a critical “Generalist Gap”: a lack of consensus on a comprehensive model for non-procedural or primary care physicians who will increasingly rely on AI-driven decision support. This paper addresses that structural gap. We move beyond the question of why AI should be taught to the practical challenge of how. By comparing existing models and synthesizing current best practices, we present the AI-PACE Framework (Psychomotor, Affective, Cognitive, Embedded).
This model runs longitudinally, introducing AI literacy in the first year of medical school and continuing through practice. Consider that 79% of U.S. medical schools begin physical exam training within the first two months of the pre-clerkship phase to establish early competency (Uchida and others, 2019). The same logic applies here: early and repeated exposure to AI concepts builds the foundation physicians need to work effectively with these tools as they advance through training. Without this preparation, physicians risk becoming passive consumers of AI outputs rather than professionals equipped to evaluate and apply them appropriately.
2 Methods
To develop a comprehensive framework for AI in medical education, we conducted a systematic search and comparative analysis of the literature. We queried scholarly databases (PubMed, MEDLINE, ERIC) for peer-reviewed articles published between 2016 and 2025, using keywords related to artificial intelligence, medical education curricula, competency frameworks, and physician training. We also used Asta (Allen AI, 2025), from Allen AI to search for additional papers. This timeframe was selected to capture the rapid evolution of clinical AI applications and the corresponding educational responses.
Selection Criteria
Our initial search yielded 643 unique records. We screened these for relevance based on three inclusion criteria:
  1. 1.
    The proposal or evaluation of a structured AI curriculum or competency framework;
  2. 2.
    A target audience of medical students (Undergraduate Medical Education: UME), residents (Graduate Medical Education: GME), or practicing physicians (Continuing Medical Education: CME);
  3. 3.
    Publications in English.
We excluded purely technical papers, commentaries without curricular proposals, and studies focused on medical training outside of the United States. This process resulted in 23 key papers identified for in-depth review.
Analytical Approach
We performed a thematic analysis of the selected literature to identify recurring competencies and structural models. Frameworks were evaluated against three critical dimensions distilled from modern medical education theory:
  • Longitudinal Integration: Does the model support learning across the continuum of training (UME, GME, CME)?
  • Generalizability: Is the framework applicable to general medical practice, or is it restricted to specific specialties (e.g., Radiology)?
  • Domain Breadth: Does the framework integrate technical knowledge (Cognitive) with workflow application (Psychomotor) and professional attitudes (Affective)?
This comparative analysis (summarized in Table 1) revealed specific gaps in existing models, namely, the fragmentation of training into isolated workshops and the under-representation of the affective domain. The AI-PACE Framework operationalizes these competencies by grounding them in the Psychomotor, Affective, and Cognitive domains of Bloom’s Taxonomy (Anderson and Krathwohl, 2001), while adding “Embedded” as a structural pillar for longitudinal delivery.
Table 1: Comparative Analysis of Existing Frameworks
Study / Framework Target Audience & Scope Methodology Longitudinal Integration? Domains Covered Identified Gaps (vs. AI-PACE)
Valikodath et al. (Valikodath and others, 2021) (2021) Ophthalmology (UME to Fellowship) Perspective / Task Force Recommendations Yes (Proposed progression from med school to fellowship) Cog: Basics, Stats
Psycho: Surgical simulation
Aff: "Humanization" mentioned
Specialty-specific; lacks a generalized implementation model for all of medicine.
van Kooten et al. (van Kooten and others, 2024) (2024) Radiology (Residency) Pilot Study / Implementation No (3-day "fast-track" intensive course) Cog: Fundamentals
Psycho: Hands-on algorithm building
Ethics: Group discussion
Short-term intervention rather than longitudinal; specialty-specific; participants noted "hands-on" coding was less useful.
Singla et al. (Singla and others, 2024) (2024) UME (General Medical Students) Delphi Study (Expert Consensus) Suggested (Integration into existing curriculum recommended) Cog: Theory, Ethics, Law
Psycho: Application
Aff: Communication, Collaboration
Focused heavily on what to teach (competencies) rather than how to structure the educational continuum (UME->GME->CME).
Turner et al. (Turner et al., 2025) (2025) PCCM (Pulmonary & Critical Care) Perspective / Conceptual Partial (Mentions "Scalable Oversight") Cog: Ethics, Robustness
Psycho: Human-AI collaboration
Focuses on principles of safe AI integration rather than educational curriculum design; specialty-focused.
Masters (Masters, 2020) (2020) General (Admin & Educators) Conceptual Framework Conceptual (Curriculum mapping discussed) Cog: AI Literacy
Admin: Curriculum design
High-level conceptual guide for administrators; less focused on the practical syllabus for learners.
AI-PACE Framework General Medicine (UME, GME, CME) Integrative Review & Educational Theory Yes (Spiral curriculum across the full continuum) Psychomotor (Workflow)
Affective (Trust)
Cognitive (Knowledge)
Embedded (Longitudinal)
Addresses the Gap: Provides a unified, generalist model that spans the entire career continuum and explicitly operationalizes the Affective domain.
3 Current State of AI Education
Despite the proliferation of AI technologies in clinical practice, our analysis reveals a significant lag in the corresponding evolution of medical education frameworks. While recent efforts have begun to define necessary competencies, the landscape remains fragmented. A scoping review by Tolentino (Tolentino et al., 2024) found only two papers describing comprehensive curriculum frameworks for AI in medicine, with one focused specifically on ophthalmology. A recent scoping review by Rincón (Rincón and others, 2025) further confirmed these findings, revealing a lack of standardized AI curriculum frameworks and notable global discrepancies in implementation.
3.1 Fragmented Implementation and Knowledge Decay
Current AI education often relies on short-term, intensive "bootcamps" or elective workshops. For example, van Kooten (van Kooten and others, 2024) demonstrated the efficacy of a 3-day AI curriculum for radiology residents, which successfully improved confidence and self-reported knowledge in the short term. However, such concentrated burst of educational modules risk knowledge decay without reinforcement. They also can fail to integrate AI literacy into the daily clinical workflow where decision-making occurs. The majority of educational interventions consist of isolated courses, workshops, or lectures rather than integrated programs (Charow and others, 2021; Grunhut et al., 2021).
3.2 Specialty Imbalance and the Generalist Gap
Existing frameworks are heavily skewed toward image-centric specialties. Where AI education does exist, it focuses on primarily on radiology and imaging (Tolentino et al., 2024). We discussed previously the efforts focused on ophthalmology, and there have been critical principles outlined for Pulmonary and Critical Care Medicine (Turner et al., 2025). While specialty-specific models offer excellent depth, AI-PACE provides the necessary transferability for general medical education. By focusing on universal competencies, such as critical appraisal of algorithms and appropriate trust calibration, the framework prepares generalist physicians to evaluate AI-driven decision support tools across any clinical domain, rather than focusing on image-specific interpretation.
3.3 The "Human" Gap: Affective Competencies
Most critically, current models frequently overlook the "human" element of human-AI interaction. Using a Delphi approach provided a strong consensus on what to teach (Singla and others, 2024), validating competencies in ethics, theory, and application, but that particular framework focuses primarily on knowledge and skills. What remains under-addressed is the Affective Domain: the calibration of trust, the recognition of automation bias, and the preservation of empathy in an AI-mediated patient encounter. Rincón highlighted an emerging educational need for strengthening transversal skills, such as ethical awareness, to support the integration of AI as a practical tool rather than a standalone subject (Rincón and others, 2025).
4 The AI-PACE Framework
To synthesize our findings and address the structural gaps in current medical education, we developed the AI-PACE Framework. This model organizes competencies into four pillars, grounding the content in the established domains of Bloom’s Taxonomy (Psychomotor, Affective, Cognitive) while introducing a fourth structural dimension (Embedded) to ensure longitudinal integration across the medical training continuum (Figure 1).
Figure 1: AI-PACE Taxonomic Framework for Medical AI Competencies within Bloom’s Taxonomy Domains.
We start by examining the mapping to Bloom’s taxonomy and then move on to the embedded approach.
4.1 Psychomotor Domain (Skills & Workflow)
The Psychomotor domain focuses on the shift from theoretical knowledge to clinical utility, emphasizing the “hands-on” ability to navigate tools, interpret probabilistic outputs, and communicate these findings to patients.
  • Operational Skills: Navigating AI tools in clinical workflows, inputting data effectively, and retrieving meaningful results (Tolentino et al., 2024).
  • Critical Appraisal: Evaluating AI algorithms for clinical validity, bias, and applicability to specific patient populations (Charow and others, 2021). This includes interpreting key performance metrics (sensitivity, specificity, AUC) without necessarily calculating them.
  • Data Analysis: Interpreting outputs from AI systems and integrating them into clinical decision-making (Charow and others, 2021).
  • Communication: Explaining AI-driven recommendations to patients in understandable terms and translating technical concepts (Masters, 2020).
4.2 Affective Domain (Appropriate Entrustment & Trust Calibration)
The affective domain encompasses the attitudes, values, and emotional intelligence required for human-AI partnership. This domain is critical for addressing the "human gap" identified in our analysis. However, framing this as "trust" undersells what clinical training actually requires. Clinicians must learn to assess AI outputs as they would a junior colleague’s work. This means understanding where the tools can fail (i.e. LMMs summarize effectively but struggle to extract qualitative meaning from patient narratives). Therefore, curricula must address these boundaries explicitly rather than treating AI competence as a binary trust decision.
  • Appropriate Trust Calibration: Developing a balanced perspective on AI capabilities to avoid both "automation bias" (over-reliance) and "algorithm aversion" (unwarranted skepticism) (Grunhut et al., 2022).
  • Patient-Centered Values: Maintaining focus on humanistic care while leveraging technological tools (Li et al., 2019). AI should augment rather than replace empathy and compassion.
  • Collaboration: Valuing interdisciplinary teamwork with data scientists and engineers (Charow and others, 2021).
  • Continuous Learning: Fostering commitment to ongoing education as AI technologies evolve (Wartman and Combs, 2018).
4.3 Cognitive Domain (Knowledge)
The cognitive domain establishes the foundational knowledge base required to understand AI as a tool.
  • Fundamentals of AI: Understanding basic concepts, terminology (machine learning, deep learning), and knowing that ML models generate probabilistic predictions based on patterns, not rigid rules (Masters, 2020; Paranjape et al., 2019). Foundational training must also update outdated skepticism. The 2022-era critique of AI as "stochastic parrots", probabilistic word repeaters, no longer reflects current model capabilities (OpenAI, 2025; Google, 2025; Anthropic, 2025). Reasoning and agentic functions have advanced substantially. Curricula should ground skepticism in actual technical limitations.
  • Health Data Science: Comprehending biostatistics, big data principles, and data governance (Charow and others, 2021).
  • Strengths and Limitations: Recognizing the capabilities and constraints of AI systems, including sources of error and clinical relevance (Grunhut et al., 2021).
  • Ethical and Legal Considerations: Understanding patient privacy, algorithmic bias, and regulatory frameworks governing AI in healthcare (Masters, 2020; Pregowska and Perkins, 2024).
4.4 Embedded (Longitudinal Integration)
The "Embedded" pillar addresses the structural limitations of fragmented "bootcamp" models by distributing AI competencies across the full medical training continuum, which aligns with the higher levels of Harden’s “integration ladder” (Harden, 2000). Rather than treating AI as an isolated elective or a standalone "AI week," this approach aligns with Han’s work (Han and others, 2019) by weaving AI literacy into the existing fabric of pre-clerkship and clerkship coursework.
This longitudinal model directly bridges the "Generalist Gap" by establishing foundational cognitive and affective competencies early in training, which then branch into specialized psychomotor skills during residency and independent practice.
Implementation must also address the faculty gap. Early-career faculty often have strong interest in AI literacy but may hesitate to enter classes where students could outpace them. This creates a disconnect: the faculty best positioned to integrate emerging tools might also be those most wary of the optics of learning along with their students. Therefore, parallel upskilling pathways for junior faculty are essential, without them, we cannot expect faculty to model the critical evaluation and entrustment behaviors the curriculum intends to teach.
  • Undergraduate Medical Education (UME): Focuses on foundational concepts and basic ethical awareness. AI-specific content is integrated into existing modules; for example, the cognitive distinction between deterministic algorithms and probabilistic machine learning is taught alongside Evidence-Based Medicine (EBM) and Biostatistics.
  • Graduate Medical Education (GME): Shifts toward specialty-specific applications and workflow integration. Residents move from theory to clinical utility, practicing "human-in-the-loop" verification during specialty rotations (e.g., Internal Medicine or Radiology clerkships) through hands-on, case-based learning.
  • Continuing Medical Education (CME): Emphasizes practice integration strategies, emerging technologies, and institutional leadership. Targeted workshops and self-directed learning address the evolving needs of practicing physicians, focusing on the selection, implementation, and oversight of AI tools within active workflows.
To operationalize this approach, Table 2 provides a curriculum map that aligns the AI-PACE domains with specific milestones and existing integration points across the medical education continuum. This map transitions from foundational theory in undergraduate years to advanced clinical oversight and leadership in continuing practice.
Table 2: AI-PACE Curriculum Map and Integration Strategy
Training Phase Primary Focus Cognitive (Knowledge) Psychomotor (Skills/Workflow) Affective (Attitudes/Trust) Integration Point & Learning Approach
Undergraduate (UME - Pre-clinical) Foundations AI Terminology & Concepts; ML vs. Deterministic logic; Probabilistic nature of AI; Data ethics & bias. Basic prompt engineering for study aids; Navigating literature search AI; Overview of AI applications. Recognizing automation bias; Valuing data privacy; Skepticism of "Black Box." Introductory courses & integrated modules in Epidemiology, Biostatistics, & Health Systems Science.
Undergraduate (UME - Clinical) Exposure Interpreting AI metrics (AUC, Sensitivity); Regulatory landscapes (FDA); Data Science Basics. Critical appraisal of AI tools during rounds; Using AI for differential diagnosis; Communicating results to patients. Calibration of trust; Empathy in AI-mediated patient encounters. Clinical Clerkships (e.g., IM, Radiology) via integrated modules and clinical exposure.
Graduate (GME - Residency) Application Specialty-Specific AI Applications; Sepsis predictors; Image triage algorithms. Clinical Integration & Decision Making; "Human-in-the-loop" verification; EHR documentation; Error reporting. Professional accountability; Collaboration with data science teams. Hands-on applications & case-based learning during Morning Reports, Grand Rounds, & QI projects.
Continuing (CME - Practice) Mastery & Leadership Emerging LLM capabilities; Quality & Performance Assessment; Liability updates. Workflow Optimization; Selecting/implementing AI tools; Leadership in AI Implementation. Fostering a culture of safety; Mentoring junior staff on AI oversight. Targeted workshops, just-in-time training, and institutional leadership roles.
This outlines the Embedded (Longitudinal) pillar of AI-PACE by mapping the Cognitive, Psychomotor, and Affective domains to specific milestones in training.
Integrated Pedagogical Methods:
  • Case-Based Learning: Use specialty-specific "AI-Fail" case studies to teach trust calibration and verification.
  • Simulation Objective Structured Clinical Examination (OSCEs): Practice explaining AI-generated diagnoses and algorithmic limitations to standardized patients.
  • Quality Improvement Projects: Encourage GME/CME learners to evaluate AI tool performance and safety within their actual clinical workflows.
5 Implementation Strategy
Successfully deploying the AI-PACE framework requires a structured approach to curriculum development that can keep pace with rapid technological shifts. Based on established models by Harden (Harden, 1986) and Thomas (Thomas et al., 2022), we put forward a six-step cycle specifically adapted for the dynamic nature of clinical AI (Figure 2).
Figure 2: A proposed approach to implementing an AI curriculum into an existing program.
  1. 1.
    Define Purpose and Need: Identify gaps in current education and articulate relevance to future practice.
  2. 2.
    Establish Learning Objectives: Define specific outcomes across the Cognitive, Psychomotor, and Affective domains (as outlined in AI-PACE).
  3. 3.
    Design Course Content: Balance theoretical foundations with practical applications, tailoring content to the educational level (UME vs GME).
  4. 4.
    Organize Content Progression: Create a logical sequence building from foundational to advanced concepts, identifying integration points within the existing curriculum.
  5. 5.
    Implement Teaching Strategies: Select appropriate methods (lectures, case-based learning, projects) and develop faculty capabilities.
  6. 6.
    Evaluate and Refine: Collect systematic feedback and measure outcomes. Crucially, this step must be continuous to update content as technology evolves.
6 Discussion
The rapid evolution of AI in medicine has outpaced the development of comprehensive educational frameworks, creating a "preparedness gap" that threatens the safe integration of these technologies into clinical practice. The AI-PACE framework addresses these limitations by offering a unified, spiral curriculum model that evolves with the learner.
6.1 The "Integration vs. Addition" Debate
A primary barrier to AI education is the "crowded curriculum" problem. Medical school and residency schedules are already saturated (Wartman, 2019). Integrating AI concepts into existing courses rather than creating new courses can help address this challenge (Grunhut et al., 2022). For example, the cognitive distinction between deterministic algorithms and probabilistic machine learning should be taught alongside EBM during pre-clerkship years. Similarly, the psychomotor skill of "human-in-the-loop" verification should be practiced during existing clinical clerkships, explicitly avoiding the “fragmented bootcamp” model that detaches technology from the daily clinical workflow. This integration model mirrors how EBM itself was once a novel addition that is now woven into the fabric of clinical reasoning.
6.2 The Resource Challenge: Faculty and Funding
Implementing AI-PACE requires overcoming significant systemic hurdles. Many medical educators lack expertise in AI and data science (Grunhut et al., 2022; Chan and Zary, 2019). This can be addressed through faculty development programs, collaboration with computer science departments, and recruitment of clinician-informaticians. Briganti and Le Moine highlight that many schools are addressing this by associating medical curricula with engineering curricula (Briganti and Le Moine, 2020). Furthermore, the fast pace of AI development risks making curricular content quickly outdated (Charow and others, 2021). Institutions must leverage interdisciplinary partnerships, breaking down the silos between Computer Science, Informatics, and Medicine, to share resources and ensure content remains current (Masters, 2020).
6.3 Validation and Future Directions
As medical education adopts AI curricula, we must rigorously validate their impact. Future research should move beyond self-reported confidence (Kirkpatrick Level 1) to measure changes in behavior (Level 3) and patient outcomes (Level 4) (Kirkpatrick and Kirkpatrick, 2006; Yardley and Dornan, 2012). Does an AI-literate resident make safer decisions when using a diagnostic support tool? By standardizing the educational model through AI-PACE, we can begin to generate the multi-institutional data needed to answer these questions. This is critical because we are learning that relying on GenAI for efficiency may come at a cost of skill formation (Shen and Tamkin, 2026).
7 Conclusion
The integration of AI into healthcare necessitates corresponding changes in medical education. Our analysis of the literature reveals a significant gap between the rapid adoption of AI in clinical practice and the preparation of physicians to use these technologies effectively. While a limited number of educational initiatives exist, comprehensive curriculum frameworks are largely absent or fragmented. We have identified key competencies across cognitive, psychomotor, and affective domains that should guide AI education for medical professionals. Effective curricula should balance technical knowledge with clinical applications and ethical considerations, be integrated longitudinally throughout medical training, and involve interdisciplinary collaboration.
The AI-PACE framework presented in this paper offers a starting point for medical educators seeking to develop AI curricula. Future research should focus on implementing and evaluating such curricula, developing specialty-specific educational approaches, and creating faculty development programs to support AI education. As AI continues to transform healthcare, ensuring that physicians are adequately prepared to leverage these technologies while maintaining humanistic, patient-centered care remains a critical challenge for medical education.